Solve not in excel and provide all the formulas

You are a sales representative of an Estonian company and your job is to broker domestic wild berries to the UK market. You have just reached an agreement to deliver 600 tonnes of berries to UK at a price of £ 5.5 per kilogram. The average purchase price for berries is EUR 5.00 per kilogram based on contracts with your suppliers. You can expect all cash flows to occur exactly one year later. Current exchange rate is 0.8400EURGBP.

1) Explain what is the direction of exchange rate risk here and how this can be mitigated using forward contract? Suppose for a moment that the interest rate in the euro area is 2.40% and in the United Kingdom comparable interest rate is 1.1 per annum 2) Find the EUR/GBP forward rate for one year based on interest parity 3) Calculate the profit per transaction (in EUR) based on the forward rate?
Answer #1

1]

The risk is that the EUR/GBP exchange rate will rise. If the EUR/GBP exchange rate rises, each GBP will realize a lower number of EUR. Since the sale proceeds are received in GBP and then converted into EUR, a rise in EUR/GBP exhange rate will result in lower number of GBP received per tonne of berries.

2]

Forward EUR/GBP rate in 1 year = spot exchange rate * (1 + UK interest rate) / (1 + Euro Interest rate)

Forward EUR/GBP rate in 1 year = 0.8400 * (1 + 1.1%) / (1 + 2.4%)

Forward EUR/GBP rate in 1 year = 0.8293

3]

Profit per kilogram = sale price (in EUR) - purchase price (in EUR)

Sale price in EUR = sale price in GBP /  Forward EUR/GBP rate

Sale price in EUR = (£5.5 / 0.8293) = 6.6321

Profit per kilogram = 6.6321 - 5.00 = EUR 1.6321

Profit on 600 tonnes - 1.6321 * 600 * 1000 = EUR 979,259.62